# The resummation model in FLAMINGO: precisely predicting matter power suppression from observed halo baryon fractions

### v1

@@ Abstract
In order to derive unbiased cosmological parameters from Stage-IV surveys, we need models that can predict the matter power spectrum for at least $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$ with percent-level precision. The main challenge in this endeavour is that baryonic feedback significantly redistributes matter on large scales, but to an unknown degree. Here, we present an improved version of the “resummation” model, which maps observed halo baryon fractions of massive haloes ($M_{\mathrm{500c}}\gtrsim 10^{12.5}\,\mathrm{M_{\odot}}$) to a flexible suppression signal – i.e. the ratio of baryonic to dark-matter-only (DMO) matter power spectra – using zero free parameters. We apply this model to the FLAMINGO hydrodynamical simulations, obtaining a typical precision of $\lesssim 1\%$ for $k\,{\leq}\,{10}\,h\mathrm{\,Mpc^{-1}}$ given mean halo baryon fractions within the spherical overdensity radii $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$. When only those within $R_{\mathrm{500c}}$ are available, we still obtain $\lesssim 2\%$ precision. We show that given small-scale stellar mass fractions, the model can be extended to yield $\lesssim 3\%$ precise suppression signals for all scales measured ($k\,{\leq}\,{25}\,h\mathrm{\,Mpc^{-1}}$). We also extend the model to redshifts $z>0$. Central to the model is a seemingly mass-independent and feedback-independent relation that allows observed halo masses to be mapped to equivalent DMO halo masses using only observed mean halo baryon fractions, to $\lesssim 1\%$ precision. We demonstrate that this relation can also be used to retrieve the DMO halo mass function from observed halo masses and baryon fractions with percent-level precision, without any assumptions on the strength of feedback. A Python package implementing the resummation model is made publicly available.
@@ 1 Introduction
One of the main goals of cosmology – to precisely constrain the parameters that determine the Universe we live in – is now more within reach than ever, thanks to nascent and upcoming Stage-IV surveys like ESA’s Euclid [^1] https://www.euclid-ec.org/, Vera C. Rubin Observatory’s LSST [^2] https://rubinobservatory.org/, the Roman wide-field survey [^3] https://roman.gsfc.nasa.gov/ and the Simons Observatory  [^4] https://simonsobservatory.org/. These surveys will measure matter clustering to unprecedented precision, which can in principle be connected to the cosmology of our Universe. In order to predict measures like the matter power spectrum with $\sim 1\%$ precision for $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$, as required to extract all available information from these observations, dark-matter-only (DMO) two-point clustering emulators based on large suites of N-body simulations have been developed. These include FrankenEmu (Heitmann et al., 2014), EuclidEmulator2 (Euclid Collaboration: Knabenhans et al., 2021), BACCO (Angulo et al., 2021), CosmicEmu (Moran et al., 2023) and CSST Emulator (Chen et al., 2025), and are continuously improved to yield higher accuracy, extend to larger wave numbers $k$, and cover more diverse cosmologies.
@@
However, as van Daalen et al. (2011) and Semboloni et al. (2011) first put forth, the main challenge in correctly interpreting measurements of the clustering of matter and coupling these to a cosmology in an unbiased manner, is that the baryonic feedback processes associated with galaxy formation redistribute matter on large scales. One often-studied measure of this is the ratio of the matter power spectrum of a realistic scenario (i.e. one including baryons) to a power spectrum predicted for a DMO Universe with the same cosmology. Since on the most relevant scales for weak lensing surveys ($k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$), baryons smooth the matter distribution and therefore suppress clustering, we will refer to this ratio $P_{\mathrm{hydro}}(k)/P_{\mathrm{DMO}}(k)$ as the “power suppression signal” – even though on smaller scales, corresponding to inner haloes and galaxies, baryons greatly enhance the power relative to DMO. While the accuracy, volumes and resolution with which galaxies are modelled in simulations have immensely improved since this systematic bias was discovered, galaxy formation is not a solved problem, and a lot of freedom remains in how processes like feedback from active galactic nuclei (AGN) redistribute gas, how much gas is ejected, and out to what scales.
@@
Fortunately, the wealth of data measuring the distributions of gas or matter as a whole, collected in recent years, has opened up many avenues for constraining the effects of baryons with observations. These now include hot gas fractions and cluster profiles measured in X-ray (e.g. Akino et al., 2022; Grandis et al., 2024; Ferreira et al., 2024; Bahar et al., 2024), the kinetic Sunyaev-Zel’dovich effect (kSZ, e.g. Hadzhiyska et al., 2024), the thermal Sunyaev-Zel’dovich effect with weak-lensing-calibrated halo masses (tSZ, e.g. To et al., 2024; Dalal et al., 2025), and, increasingly, combinations of multiple such probes, often combined with cosmic shear, galaxy-galaxy or CMB lensing (e.g. Tröster et al., 2022; Schneider et al., 2022; Chen et al., 2023; Aricò et al., 2023; García-García et al., 2024; Bigwood et al., 2024; McCarthy et al., 2025; Sunseri et al., 2025; Kovač et al., 2025; Hadzhiyska et al., 2025). Since very recently, it has even become possible to probe the effects of feedback through the dispersion measure as probed by fast radio bursts (FRBs, e.g. Leung et al., 2025; Reischke & Hagstotz, 2025), as first proposed by Reischke et al. (2023). All these constraints, certainly when used in combination, have already ruled out a number of models. Despite this, the precise magnitude of baryonic effects in our Universe is still highly uncertain, especially on small scales, and can differ a lot per study, with results depending on the type of observable, the instrument, and the method used to interpret the data. To get the most out of large-scale clustering surveys, we therefore need both more observations and more precise models.
@@
Central to these latter efforts are hydrodynamical simulations, particularly suites that explore a range in feedback strengths and implementations in relevant volumes. These include cosmo-OWLS (Le Brun et al., 2014), BAHAMAS (McCarthy et al., 2017), CAMELS (Villaescusa-Navarro et al., 2021), ANTILLES (Salcido et al., 2023), MillenniumTNG (Hernández-Aguayo et al., 2023), FLAMINGO (Schaye et al., 2023; Kugel et al., 2023), FABLE and XFABLE (Martin-Alvarez et al., 2025; Bigwood et al., 2025). In the context of matter clustering, such simulations can not only be used to quantify the effects of feedback, but also to test modelling and mitigation strategies for the effects of baryons, and, for example, develop measures for the redistribution of gas on large scales (e.g. Gebhardt et al., 2024; Medlock et al., 2025; Sharma et al., 2025). As argued by van Daalen et al. (2020), one such very informative quantity is the mean halo baryon fraction, particularly on the group scale ($\sim 10^{14}\,\mathrm{M_{\odot}}$). This quantity can by itself accurately predict the suppression signal across a wide range of feedback strengths and implementations, at least on scales $k\,{\lesssim}\,{1}\,h\mathrm{\,Mpc^{-1}}$. It is the baryon fraction – i.e. the mass in gas and stars within some radius around halo centres relative to the total mass therein – rather than just the gas fraction, that reveals how much baryonic matter has been driven out of the halo, and redistributed to larger scales. After all, a scenario in which massive haloes overcool and lock much of their gas content in stars, and another in which AGN feedback is highly effective, may have similar gas fractions, but will have very different baryon fractions, and very different large-scale matter distributions. Only at fixed stellar fraction, then, are gas fractions a measure of feedback. Observations of halo gas profiles and gas content are therefore best combined with optical and infrared surveys measuring the stellar content of haloes whenever possible, such as done in e.g. Akino et al. (2022) and Dev et al. (2024).
@@
Many modelling efforts and mitigation strategies concerning the effects of baryons on clustering, based on the results of hydrodynamical simulations, currently exist (e.g. Semboloni et al., 2011, 2013; Zentner et al., 2013; Huang et al., 2021; Salcido et al., 2023; Schaller et al., 2025; Kammerer et al., 2025), including the popular hmcode (Mead et al., 2015, 2021), the highly successful baryonification/baryon-correction method approaches (e.g. Schneider & Teyssier, 2015; Aricò et al., 2020, 2021; Schneider et al., 2019, 2025), and direct modifications of the power spectral shape (e.g. Amon & Efstathiou, 2022; Preston et al., 2023; Schaller & Schaye, 2025). However, while these are of great use, even when based on the results of hydrodynamical simulations, all of these methods need to either assume limited functional forms for density profiles or the suppression signal itself, have free parameters that need to be marginalised over, or both. This often limits their flexibility, i.e. their ability to model suppression signals they were not designed for or calibrated to, and/or increases the uncertainties in the derived set of cosmological parameters. The former is an issue as the theory of galaxy formation is still being expanded with new processes that may play an active role in redistributing gas in ways that were not previously modelled. Examples include bipolar jets (e.g. Davé et al., 2019; Schaye et al., 2023) and cosmic ray feedback, which Quataert & Hopkins (2025) recently argued may have a particularly large effect on the gas distribution around and beyond the virial radii of groups.
@@
With all of this in mind, it is worthwhile to consider whether the effects that baryons have on haloes and the matter distribution at large scales can be modelled more agnostically, while simultaneously keeping the number of free parameters to a minimum. Recently, van Loon & van Daalen (2024) investigated the role of haloes of different masses in shaping the matter power spectrum, with and without feedback, building on the work of Semboloni et al. (2011), Semboloni et al. (2013) and van Daalen & Schaye (2015). Their findings demonstrated why the baryon fractions of group-size haloes of mass $M_{\mathrm{500c}}\approx 10^{14}\,\mathrm{M_{\odot}}$ is strongly correlated with the suppression of power on large scales: these haloes have about the optimal combination of mass and halo bias to provide the dominant contribution to the total power on scales $k\,{\lesssim}\,{2}\,h\mathrm{\,Mpc^{-1}}$ (a result that was further explored by Salcido et al., 2023). Additionally, the baryon fraction in haloes of adjacent mass scales, which provide the second-most dominant contribution, can reasonably be expected to correlate with the baryon fractions of the $10^{14}\,\mathrm{M_{\odot}}$ haloes themselves. This offered an explanation for the strong correlation between power suppression and baryon fractions of massive groups found by van Daalen et al. (2020), though it also pointed to some of its limits. Firstly, on smaller scales ($k\,{\gtrsim}\,{1}\,h\mathrm{\,Mpc^{-1}}$), contributions of less massive haloes and those from matter inside massive haloes all play an increasingly strong role in shaping the power spectrum, and so the correlation is weakened. Secondly, if feedback operates in such a way that the baryon fractions at, say, haloes with mass $10^{13}\,\mathrm{M_{\odot}}$ and $10^{15}\,\mathrm{M_{\odot}}$ are not as tightly correlated with those of $10^{14}\,\mathrm{M_{\odot}}$ haloes, for example because AGN rapidly transition to a different feedback mode as a function of mass, then the correlation with the power suppression breaks down on even larger scales.
@@
Because of this, van Loon & van Daalen (2024) proposed a new method, dubbed the “resummation model”. The basis of the model is simple: since we can decompose the matter power spectrum into a sum of haloes cross-correlated with all matter, and a cross-term for the remaining matter not in haloes, knowing how the mass fraction in each component changes should tell us how the total signal changes – provided the shape of the redistribution beyond haloes is unimportant. In this paper, with the help of the FLAMINGO simulations, we develop the proof-of-concept version of the resummation method into a fully-fledged model, capable of predicting the power suppression signal to percent-level precision on all relevant Fourier scales $k$ given observed halo baryon fractions – without free parameters, and without assuming a profile for the matter distribution or power spectral shape.
@@
This paper is structured as follows. In Section 2, we introduce the FLAMINGO simulations and the resummation model, including several improvements and extensions relative to van Loon & van Daalen (2024). Then, in Section 3, we demonstrate the seemingly-universal relation between the mass ratio of a hydrodynamical halo and its DMO equivalent as a function of its baryon fraction, regardless of its mass, cosmology or feedback implementation. This relation, presented in §3.1, is one of our main results. Then, continuing Section 3, we test the model’s ability to predict the power suppression signal of all FLAMINGO simulations, given various levels of information about the mean baryon fractions of haloes, showing it achieves percent-level precision. Next, in Section 4, we discuss our findings and the strengths and limitations of the model, predict the suppression signal of our Universe given observed baryon fractions within $R_{\mathrm{500c}}$, and show that it can also be used to accurately recover the DMO halo mass function without additional assumptions or parameters, given those same baryon fractions. Finally, we summarise our work in Section 5.
@@
A Python package, resummation, along with the necessary DMO cross spectra and other quantities for the FLAMINGO simulations, will be made publicly available upon acceptance of this manuscript for publication.
@@ 2.1 Simulations
The FLAMINGO simulations, described in detail in Schaye et al. (2023) and Kugel et al. (2023), are a suite of large-volume, cosmological hydrodynamical simulations, calibrated to simultaneously reproduce both the $z=0$ galaxy stellar mass function (SMF) and gas fractions ($f_{\mathrm{gas}}$) of massive groups and clusters in the redshift range z=0.10.3z=0.1-0.3.555We note that the gas fractions used to calibrate FLAMINGO were based on pre-eROSITA X-ray and weak-lensing data, and that more recent X-ray as well as kSZ data suggest lower gas fractions are required. We return to this point in §4.2. The simulations were performed with the swift code (Schaller et al., 2024), using the SPHENIX formulation (Borrow et al., 2022) of Smoothed Particle Hydrodynamics (SPH) and the $\delta f$ method for evolving neutrinos of Elbers et al. (2021). The initial conditions (ICs) were generated using the MonofonIC code (Hahn et al., 2021; Elbers et al., 2022) using a 3-fluid formalism with a separate transfer function or each of the dark matter, gas, and neutrinos. The ICs used partially fixed modes (Angulo & Pontzen, 2016), setting the amplitudes of modes with $(kL)^{2}<1025$ to the mean variance, where $L$ is the side length of the simulated volume and $k$ is the wave number of a mode. The multi-fluid approach (Rampf et al., 2021) used by MonofocIC allows for a perfect match of the gravity-only and full-hydrodynamics power spectrum on the largest scales.
@@
The subgrid prescriptions of FLAMINGO include element-by-element radiative cooling and heating (Ploeckinger & Schaye, 2020), star formation (Schaye & Dalla Vecchia, 2008), stellar mass loss (Wiersma et al., 2009; Schaye et al., 2015), stellar feedback (Chaikin et al., 2022, 2023), and the growth of supermassive black holes (Springel et al., 2005; Booth & Schaye, 2009; Bahé et al., 2022). Feedback from active galactic nuclei (AGN) is implemented either thermally following Booth & Schaye (2009), which is the case for most simulations, or following the kinetic jet-based prescription of Huško et al. (2022), which is the case for the two simulations with “Jet” in their name.
@@
The simulation suite has several properties beneficial to our investigation. Firstly, they vary the parameters governing stellar and AGN feedback in such a way as to reproduce the SMF and/or $f_{\mathrm{gas}}$ data shifted up or down by $\pm N\sigma$ – where $\sigma$ is the relevant error on the observational data – following the procedure detailed in Kugel et al. (2023). Secondly, they vary the cosmological parameters of the simulations relative to the fiducial “3x2pt + all external constraints” Dark Energy Survey year 3 results of Abbott et al. (2022). Thirdly, every hydrodynamical FLAMINGO simulation has a dark-matter-only (DMO) counterpart, with the same cosmology and initial conditions. By considering simulations with different feedback implementations and cosmologies, and comparing their matter power spectra to that of their DMO counterparts, we can investigate a range of baryonic suppression signals, and explore the performance of the resummation model in varying scenarios.

### v2

@@ Abstract
In order to derive unbiased cosmological parameters from Stage-IV surveys, we need models that can predict the matter power spectrum for at least $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$ with percent-level accuracy. The main challenge in this endeavour is that baryonic feedback significantly redistributes matter on large scales, but to an unknown degree. Here, we present an improved version of the “resummation” model, which maps observed halo baryon fractions of massive haloes ($M_{\mathrm{500c}}\gtrsim 10^{12.5}\,\mathrm{M_{\odot}}$) to a flexible suppression signal – i.e. the ratio of baryonic to dark-matter-only (DMO) matter power spectra – using zero free parameters. We calibrate this model to the FLAMINGO hydrodynamical simulations, obtaining a typical accuracy of $\lesssim 1\%$ for $k\,{\leq}\,{10}\,h\mathrm{\,Mpc^{-1}}$ given mean halo baryon fractions within the spherical overdensity radii $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$. When only those within $R_{\mathrm{500c}}$ are available, we still obtain $\lesssim 2\%$ accuracy. We show that given small-scale stellar mass fractions, the model can be extended to yield $\lesssim 3\%$ accurate suppression signals for all scales measured ($k\,{\leq}\,{25}\,h\mathrm{\,Mpc^{-1}}$). We also extend the model to redshifts $z>0$. Central to the model is a seemingly mass-independent and feedback-independent relation that allows observed halo masses to be mapped to equivalent DMO halo masses using only observed mean halo baryon fractions, to $\lesssim 1\%$ accuracy. This relation can also be used to retrieve the DMO halo mass function from observed halo masses and baryon fractions with percent-level accuracy, without any assumptions on the strength of feedback. A Python package implementing the resummation model is made publicly available.
@@ 1 Introduction
One of the main goals of cosmology – to precisely constrain the parameters that determine the Universe we live in – is now more within reach than ever, thanks to nascent and upcoming Stage-IV surveys like ESA’s Euclid [^1] https://www.euclid-ec.org/, Vera C. Rubin Observatory’s LSST [^2] https://rubinobservatory.org/, the Roman wide-field survey [^3] https://roman.gsfc.nasa.gov/ and the Simons Observatory  [^4] https://simonsobservatory.org/. These surveys will measure matter clustering to unprecedented precision, which can in principle be connected to the cosmology of our Universe. In order to predict measures like the matter power spectrum with $\sim 1\%$ precision for $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$, as required to extract all available information from these observations, dark-matter-only (DMO) two-point clustering emulators based on large suites of N-body simulations have been developed. These include FrankenEmu (Heitmann2014), EuclidEmulator2 (Euclid2021), BACCO (Angulo2021), CosmicEmu (Moran2023) and CSST Emulator (Chen2025), and are continuously improved to yield higher accuracy, extend to larger wave numbers $k$, and cover more diverse cosmologies.
@@
However, as vanDaalen2011 and Semboloni2011 first put forth, the main challenge in correctly interpreting measurements of the clustering of matter and coupling these to a cosmology in an unbiased manner, is that the baryonic feedback processes associated with galaxy formation redistribute matter on large scales. One often-studied measure of this is the ratio of the matter power spectrum of a realistic scenario (i.e. one including baryons) to a power spectrum predicted for a DMO Universe with the same cosmology. Since on the most relevant scales for weak lensing surveys ($k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$), baryons smooth the matter distribution and therefore suppress clustering, we will refer to this ratio $P_{\mathrm{hydro}}(k)/P_{\mathrm{DMO}}(k)$ as the “power suppression signal” – even though on smaller scales, corresponding to inner haloes and galaxies, baryons greatly enhance the power relative to DMO. While the accuracy, volumes and resolution with which galaxies are modelled in simulations have immensely improved since this systematic bias was discovered, galaxy formation is not a solved problem, and a lot of freedom remains in how processes like feedback from active galactic nuclei (AGN) redistribute gas, how much gas is ejected, and out to what scales.
@@
Fortunately, the wealth of data measuring the distributions of gas or matter as a whole, collected in recent years, has opened up many avenues for constraining the effects of baryons with observations. These now include hot gas fractions and cluster profiles measured in X-ray (e.g. Akino2022; Grandis2024; Ferreira2024; Bahar2024), the kinetic Sunyaev-Zel’dovich effect (kSZ, e.g. Hadzhiyska2025a), the thermal Sunyaev-Zel’dovich effect with weak-lensing-calibrated halo masses (tSZ, e.g. To2024; Dalal2025), and, increasingly, combinations of multiple such probes, often combined with cosmic shear, galaxy-galaxy or CMB lensing (e.g. Troster2022; Schneider2022; Chen2023; Arico2023; Garcia-Garcia2024; Bigwood2024; McCarthy2025; Sunseri2025; Kovac2025; Hadzhiyska2025b). Since very recently, it has even become possible to probe the effects of feedback through the dispersion measure as probed by fast radio bursts (FRBs, e.g. Leung2025; ReischkeHagstotz2025), as first proposed by Reischke2023. All these constraints, certainly when used in combination, have already ruled out a number of models. Despite this, the precise magnitude of baryonic effects in our Universe is still highly uncertain, especially on small scales, and can differ a lot per study, with results depending on the type of observable, the instrument, and the method used to interpret the data. To get the most out of large-scale clustering surveys, we therefore need both more observations and more precise models.
@@
Central to these latter efforts are hydrodynamical simulations, particularly suites that explore a range in feedback strengths and implementations in relevant volumes. These include cosmo-OWLS (LeBrun2014), BAHAMAS (McCarthy2017), CAMELS (Villaescusa-Navarro2021), ANTILLES (Salcido2023), MillenniumTNG (Hernandez-Aguayo2023), FLAMINGO (Schaye2023; Kugel2023), FABLE and XFABLE (Martin-Alvarez2025; Bigwood2025). In the context of matter clustering, such simulations can not only be used to quantify the effects of feedback, but also to test modelling and mitigation strategies for the effects of baryons, and, for example, develop measures for the redistribution of gas on large scales (e.g. Gebhardt2024; Medlock2025b; Sharma2025). As argued by vanDaalen2020, one such very informative quantity is the mean halo baryon fraction, particularly on the group scale ($\sim 10^{14}\,\mathrm{M_{\odot}}$). This quantity can by itself accurately predict the suppression signal across a wide range of feedback strengths and implementations, at least on scales $k\,{\lesssim}\,{1}\,h\mathrm{\,Mpc^{-1}}$. It is the baryon fraction – i.e. the mass in gas and stars within some radius around halo centres relative to the total mass therein – rather than just the gas fraction, that reveals how much baryonic matter has been driven out of the halo, and redistributed to larger scales. After all, a scenario in which massive haloes overcool and lock much of their gas content in stars, and another in which AGN feedback is highly effective, may have similar gas fractions, but will have very different baryon fractions, and very different large-scale matter distributions. Only at fixed stellar fraction, then, are gas fractions a measure of feedback. Observations of halo gas profiles and gas content are therefore best combined with optical and infrared surveys measuring the stellar content of haloes whenever possible, such as done in e.g. Akino2022 and Dev2024.
@@
Many modelling efforts and mitigation strategies concerning the effects of baryons on clustering, based on the results of hydrodynamical simulations, currently exist (e.g. Semboloni2011; Semboloni2013; Zentner2013; Huang2021; Salcido2023; Schaller2025; Kammerer2025), including the popular hmcode (Mead2015; Mead2021), the highly successful baryonification/baryon-correction method approaches (e.g. SchneiderTeyssier2015; Arico2020; Arico2021; Schneider2019; Schneider2025), and direct modifications of the power spectral shape (e.g. AmonEfstathiou2022; Preston2023; SchallerSchaye2025). However, while these are of great use, even when based on the results of hydrodynamical simulations, all of these methods need to either assume limited functional forms for density profiles or the suppression signal itself, have free parameters that need to be marginalised over, or both. This often limits their flexibility, i.e. their ability to model suppression signals they were not designed for or calibrated to, and/or increases the uncertainties in the derived set of cosmological parameters. The former is an issue as the theory of galaxy formation is still being expanded with new processes that may play an active role in redistributing gas in ways that were not previously modelled. Examples include bipolar jets (e.g. Dave2019; Schaye2023) and cosmic ray feedback, which QuataertHopkins2025 recently argued may have a particularly large effect on the gas distribution around and beyond the virial radii of groups.
@@
With all of this in mind, it is worthwhile to consider whether the effects that baryons have on haloes and the matter distribution at large scales can be modelled more agnostically, while simultaneously keeping the number of free parameters to a minimum. Recently, vanLoonvanDaalen2024 investigated the role of haloes of different masses in shaping the matter power spectrum, with and without feedback, building on the work of Semboloni2011, Semboloni2013 and vanDaalenSchaye2015. Their findings demonstrated why the baryon fractions of group-size haloes of mass $M_{\mathrm{500c}}\approx 10^{14}\,\mathrm{M_{\odot}}$ is strongly correlated with the suppression of power on large scales: these haloes have about the optimal combination of mass and halo bias to provide the dominant contribution to the total power on scales $k\,{\lesssim}\,{2}\,h\mathrm{\,Mpc^{-1}}$ (a result that was further explored by Salcido2023). Additionally, the baryon fraction in haloes of adjacent mass scales, which provide the second-most dominant contribution, can reasonably be expected to correlate with the baryon fractions of the $10^{14}\,\mathrm{M_{\odot}}$ haloes themselves. This offered an explanation for the strong correlation between power suppression and baryon fractions of massive groups found by vanDaalen2020, though it also pointed to some of its limits. Firstly, on smaller scales ($k\,{\gtrsim}\,{1}\,h\mathrm{\,Mpc^{-1}}$), contributions of less massive haloes and those from matter inside massive haloes all play an increasingly strong role in shaping the power spectrum, and so the correlation is weakened. Secondly, if feedback operates in such a way that the baryon fractions at, say, haloes with mass $10^{13}\,\mathrm{M_{\odot}}$ and $10^{15}\,\mathrm{M_{\odot}}$ are not as tightly correlated with those of $10^{14}\,\mathrm{M_{\odot}}$ haloes, for example because AGN rapidly transition to a different feedback mode as a function of mass, then the correlation with the power suppression breaks down on even larger scales.
@@
Because of this, vanLoonvanDaalen2024 proposed a new method, dubbed the “resummation model”. The basis of the model is simple: since we can decompose the matter power spectrum into a sum of haloes cross-correlated with all matter, and a cross-term for the remaining matter not in haloes, knowing how the mass fraction in each component changes should tell us how the total signal changes – provided the shape of the redistribution beyond haloes is unimportant. In this paper, with the help of the FLAMINGO simulations, we develop the proof-of-concept version of the resummation method into a fully-fledged model, capable of predicting the power suppression signal to percent-level accuracy on all relevant Fourier scales $k$ given observed halo baryon fractions – without free parameters, and without assuming a profile for the matter distribution or power spectral shape.
@@
This paper is structured as follows. In Section 2, we introduce the FLAMINGO simulations and the resummation model, including several improvements and extensions relative to vanLoonvanDaalen2024. Then, in Section 3, we demonstrate the seemingly-universal relation between the mass ratio of a hydrodynamical halo and its DMO equivalent as a function of its baryon fraction, regardless of its mass, cosmology or feedback implementation. This relation, presented in §3.1, is one of our main results. Then, continuing Section 3, we test the model’s ability to predict the power suppression signal of all FLAMINGO simulations, given various levels of information about the mean baryon fractions of haloes, showing it achieves percent-level accuracy. Next, in Section 4, we discuss our findings and the strengths and limitations of the model, predict the suppression signal of our Universe given observed baryon fractions within $R_{\mathrm{500c}}$, and show that it can also be used to accurately recover the DMO halo mass function without additional assumptions or parameters, given those same baryon fractions. Finally, we summarise our work in Section 5.
@@
A Python package, resummation, along with the necessary DMO cross spectra and other quantities for the FLAMINGO simulations, is made publicly available.
@@ 2.1 Simulations
The FLAMINGO simulations, described in detail in Schaye2023 and Kugel2023, are a suite of large-volume, cosmological hydrodynamical simulations, calibrated to simultaneously reproduce both the $z=0$ galaxy stellar mass function (SMF) and gas fractions ($f_{\mathrm{gas}}$) of massive groups and clusters in the redshift range z=0.10.3z=0.1-0.3.555We note that the gas fractions used to calibrate FLAMINGO were based on pre-eROSITA X-ray and weak-lensing data, and that more recent X-ray as well as kSZ data suggest lower gas fractions are required. We return to this point in §4.2. The simulations were performed with the swift code (Schaller2024), using the SPHENIX formulation (Borrow2022) of Smoothed Particle Hydrodynamics (SPH) and the $\delta f$ method for evolving neutrinos of Elbers2021. The initial conditions (ICs) were generated using the MonofonIC code (Hahn2021; Elbers2022) using a 3-fluid formalism with a separate transfer function or each of the dark matter, gas, and neutrinos. The ICs used partially fixed modes (AnguloPontzen2016), setting the amplitudes of modes with $(kL)^{2}<1025$ to the mean variance, where $L$ is the side length of the simulated volume and $k$ is the wave number of a mode. The multi-fluid approach (Rampf2021) used by MonofonIC allows for a perfect match of the gravity-only and full-hydrodynamics power spectrum on the largest scales.
@@
The subgrid prescriptions of FLAMINGO include element-by-element radiative cooling and heating (PloeckingerSchaye2020), star formation (SchayeDallaVecchia2008), stellar mass loss (Wiersma2009b; Schaye2015), stellar feedback (Chaikin2022; Chaikin2023), and the growth of supermassive black holes (SpringelDiMatteoHernquist2005; BoothSchaye2009; Bahe2022). Feedback from active galactic nuclei (AGN) is implemented either thermally following BoothSchaye2009, which is the case for most simulations, or following the kinetic jet-based prescription of Husko2022, which is the case for the two simulations with “Jet” in their name.
@@
The simulation suite has several properties beneficial to our investigation. Firstly, they vary the parameters governing stellar and AGN feedback in such a way as to reproduce the SMF and/or $f_{\mathrm{gas}}$ data shifted up or down by $\pm N\sigma$ – where $\sigma$ is the relevant error on the observational data – following the procedure detailed in Kugel2023. Secondly, they vary the cosmological parameters of the simulations relative to the fiducial “3x2pt + all external constraints” Dark Energy Survey year 3 results of DES2022. Thirdly, every hydrodynamical FLAMINGO simulation has a dark-matter-only (DMO) counterpart, with the same cosmology and initial conditions. By considering simulations with different feedback implementations and cosmologies, and comparing their matter power spectra to that of their DMO counterparts, we can investigate a range of baryonic suppression signals, and explore the performance of the resummation model in varying scenarios.